OpenStudy (anonymous):

Aldo's gas tank is 1/6 full. After he buys 14 gallons of gas, it is 3/4 full. How many gallons can Aldo's tank hold?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

I really don't know... Taking a guess, 1/6+14?

OpenStudy (skullpatrol):

Decide what unknown number is asked for and what facts are known.

OpenStudy (skullpatrol):

How many gallons can Aldo's tank hold? Is what is being asked for.

OpenStudy (skullpatrol):

Fact #1: Aldo's gas tank is 1/6 full. Fact #2: After he buys 14 gallons of gas, it is 3/4 full.

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so x=1/6+14=3/4 ?

OpenStudy (skullpatrol):

How much of the tank is from 1/6 full to 3/4 full?

OpenStudy (anonymous):

7/12 ?

OpenStudy (anonymous):

if i did that right, i subtracted. sorry, lol. I didn't have a lesson, it was a straight test.

OpenStudy (skullpatrol):

You did it correctly 3/4 - 1/6 = 7/12 full

OpenStudy (skullpatrol):

So now you know that (7/12) ths of the tank is how many gallons of gas?

OpenStudy (skullpatrol):

Look at fact #2

OpenStudy (anonymous):

i don't know how to figure out what that is in terms of gallons of the gas.

OpenStudy (skullpatrol):

Fact #2: After he buys 14 gallons of gas, it is 3/4 full.

OpenStudy (skullpatrol):

1/6------14gal-------->3/4

OpenStudy (anonymous):

mhm.....

OpenStudy (anonymous):

that 14 isn't the 7/12 though

OpenStudy (anonymous):

?*

OpenStudy (skullpatrol):

Sure it is.

OpenStudy (anonymous):

so, add them all up?

OpenStudy (skullpatrol):

It took 14 gals to fill 7/12ths of the tank, from 1/6 to 3/4. That is the difference between 3/4 and 1/6, so that is how much gas is between them.

OpenStudy (anonymous):

Something about that isn't right.... I can't enter fractions as an answer and the decimal is 0.583. We only figured out what was between the two.. not how many gallons total his car can hold

OpenStudy (skullpatrol):

True, that is the next step.

OpenStudy (skullpatrol):

We need to reread the problem and write an equation that represents relationships among the numbers in the problem.

OpenStudy (anonymous):

i got it. thanks

OpenStudy (skullpatrol):

Set up a proportion: $$\frac{14}{\frac{7}{12}} =\frac{x}{1}$$

OpenStudy (skullpatrol):

Thanks for trying to learn something new :)